OpenStudy (anonymous):

a car leaves an airport and drives due west at 240kph. five hours later, a plane takes off from the airport flying due west at 720kph.how far from the airport will the car be when the airplane catches it?

OpenStudy (xishem):

\[(240kph)(t)=(720kph)(t-5)\]\[240t=720t-3600\]\[-480t=-3600\]\[t=7.5hr\]\[\frac{240km}{1hr}*7.5hr=1800km\]

OpenStudy (anonymous):

thanks